# To Shuffle or not to Shuffle: Auditing DP-SGD with Shuffling

### v2

@@ 3.1 The Distinguishing Game
In Figure 2, we formalize the distinguishing game, involving an Adversary and a Challenger, used to audit the zero-out DP guarantees of a mechanism $\mathcal{M}$. The Challenger runs the mechanism on a randomly chosen dataset, and the Adversary aims to determine the dataset used by the Challenger based on the mechanism’s output.
In each game, the Adversary picks a single target record $z_{T}$ from the data domain (we slightly abuse notation and define $\mathcal{D}$ as both the data distribution and domain) and sends it to the Challenger. This constructs adjacent datasets $D_{0}$ and $D_{1}$ by appending the zero-out record $\bot$ and the target record $z_{T}$, respectively, to a randomly sampled dataset with $n-1$ records. In prior work, stronger adversarial settings where the Adversary also chooses adjacent datasets $D_{0}$ and $D_{1}$ directly have also been considered [44, 4]. However, in the context of DP-SGD, adversarial datasets can destroy model utility and represent an unrealistic adversarial setting [44, 43]. Therefore, in our work, we consider randomly crafted datasets to estimate the empirical privacy leakage in more real-world settings [43, 52, 2, 48]. Thus, the Adversary is given access to the randomly sampled adjacent datasets $D_{0}$ and $D_{1}$.
Next, the Challenger runs $\mathcal{M}$ on dataset $D_{b}$ for a random $b\in\{0,1\}$. Note that $\mathcal{M}$ can be a complex algorithm with many implementation details that the theoretical privacy analysis does not rely on; thus, depending on the threat model considered and the adversarial capabilities, the adversary may be given access to specific internals of $\mathcal{M}$ as discussed in Section 5.1, along with $M$’s output. In theory, this can take any form, e.g., scalar $o\in\mathbb{R}$, vector $o\in\mathbb{R}^{d}$, or even have arbitrary domain $o\in\mathcal{Y}$. However, it can be difficult to consider and design decision functions around values in arbitrary domains. Hence, a common strategy [44, 32] is for the Adversary to define a distinguishing function $\mathcal{A}$ that assigns a scalar “score” to the output representing the Adversary’s confidence that the output is drawn from processing $D_{1}$. This score is then thresholded to produce a guess $\hat{b}=1$ if the Adversary determines that $o\sim\mathcal{M}(D_{1})$ and 00 otherwise. The adversary wins the game if $\hat{b}=b$.
@@ 3.2 Estimating $\varepsilon_{emp}$
We estimate the empirical privacy leakage $\varepsilon_{emp}$ by running the distinguishing game multiple times, where only a single record is inserted in each run. (This is different from recent “auditing in one run” methods [52, 42, 56], which we do not use as these do not yield sufficiently tight audits of DP-SGD even under powerful threat models.) Specifically, we compute the false positive rate (FPR) $\alpha$ and the false negative rate (FNR) $\beta$ across multiple games and derive (statistically valid) upper bounds $\overline{\alpha}$ and $\overline{\beta}$ using Clopper-Pearson confidence intervals (CIs) to quantify the confidence in our privacy loss estimation [43, 32]. We then calculate the empirical lower-bound privacy guarantee $\varepsilon_{emp}$ from $\overline{\alpha}$ and $\overline{\beta}$ using the ($\varepsilon,\delta$)-DP definition as follows.
Although recent work [43, 42, 56] has proposed more advanced auditing techniques for DP-SGD (Poisson) using $f$-DP or PLD, since tighter $f$-DP guarantees for DP-SGD (Shuffle) are currently unknown [16], we stick to the ($\varepsilon,\delta$)-DP definition.
@@
$\small\begin{split}\mathcal{R}(\varepsilon,\delta)=\{(\alpha,\beta)|&\alpha+e^{\varepsilon}\beta\geq 1-\delta\land e^{\varepsilon}\alpha+\beta\geq 1-\delta~{}\land\\ &\alpha+e^{\varepsilon}\beta\leq e^{\varepsilon}+\delta\land e^{\varepsilon}\alpha+\beta\leq e^{\varepsilon}+\delta\}\end{split}$
@@
While Bayesian intervals [61] can improve the privacy estimation’s tightness, they are also more computationally expensive to derive and may not always be statistically sound [43]. Thus, we stick to Clopper-Pearson CIs. (Later in the paper, we will also evaluate the impact of using CIs on the empirical privacy leakage estimation).
@@ 3.3 The Distinguishing Function
As discussed above, DP auditing involves an adversary distinguishing between observations from $\mathcal{M}(D)$ and $\mathcal{M}(D^{\prime})$. For a mechanism $\mathcal{M}:\mathcal{X}\rightarrow\mathcal{Y}$, this can be any function of the form $\mathcal{A}:\mathcal{Y}\rightarrow\{0,1\}$ where $\mathcal{A}(o)=1$ (or 0) represents that the adversary predicts that the observation $o$ is drawn from $\mathcal{M}(D)$ (or $\mathcal{M}(D^{\prime})$).
Scalar Score. In theory, $\mathcal{A}$ is equivalent to the rejection rule from the definition of $f$-DP, which represents the adversary performing a hypothesis test on whether to reject that $o\sim\mathcal{M}(D^{\prime})$. However, in previous work [33, 32, 44, 43], the adversary assigns a scalar “score” to each output, which is then thresholded to form the distinguishing function—e.g., Jagielski et al. [32] and Nasr et al. [43] use the loss function and dot product, respectively. This is because the raw outputs of DP-SGD tend to be high-dimensional vectors where distances can be difficult to interpret, thus making distinguishing functions hard to design.
@@
Choosing the threshold. When auditing using scalar scores, choosing the threshold $\tau$ from an independent set of observations is the easiest method to ensure a technically valid lower bound for the empirical privacy leakage estimate $\varepsilon_{emp}$ derived this way. However, it is standard to report the maximum $\varepsilon_{emp}$ for the optimal threshold [3, 11, 32, 44, 43, 48, 61] and we do so too, making the process more efficient by first sorting the scalar scores. For simplicity, we abstract the details of choosing an appropriate threshold and designing a distinguishing function. We refer to the empirical estimation procedure at significance level $\alpha$ and privacy parameter $\delta$ from scores $\mathcal{S}$ and $\mathcal{S}^{\prime}$ derived from $\mathcal{M}(D)$ and $\mathcal{M}(D^{\prime})$, respectively as $\text{EstimateEps}(\mathcal{S},\mathcal{S}^{\prime},\alpha,\delta)$ – see Appendix B for more details.
4 Warm Up: Auditing the Batched Gaussian Mechanism
@@
Before auditing DP-SGD, we start with the Batched Gaussian Mechanism (BGM) (reviewed in Algorithm 4 in Appendix A), a heavily simplified version of DP-SGD adapted from [16] to develop principled tight auditing techniques for shuffling under an idealized setting. Similar to DP-SGD, BGM also proceeds iteratively, sampling batches according to some batch sampler $\mathcal{B}$. However, instead of calculating gradients, the inputs are simply aggregated together with noise at each batch, and the noisy aggregated values are released for each batch across all epochs. Also, unlike DP-SGD, it is non-adaptive, i.e., the outputs of previous iterations do not affect the current one.
@@
Although BGM can be instantiated with any batch sampler, we use the shuffle batch sampler, which is commonly used in ML training. We bound the inputs, i.e. $\forall i\;x_{i}\in[-1,+1]$, so that the mechanism satisfies $(\varepsilon,\delta)$-DP. When the number of epochs is 1, Chua et al. [16] use the adjacent datasets $D=(+1,-1,...,-1)$ and $D^{\prime}=(\bot,-1,...,-1)$, setting =0$\bot$=0⊥ = 0 in their lower bound privacy analysis. Specifically, they conjecture, but do not prove, that these are the worst-case adjacent datasets for the shuffle batch sampler. Note that this is different from the proven worst-case datasets $D=(+1,0,...,0)$ and $D^{\prime}=(\bot,0,...,0)$ for the Poisson batch sampler [62]. Specifically, in the shuffle setting, the differing sample in the conjectured worst-case datasets is equal in magnitude but opposite in direction to the other samples unlike in the Poisson setting. Finally, in a single epoch and for batch size $B$, the outputs of the mechanism on the adjacent datasets are $(-B,...,-B+2,...,-B)+\mathcal{N}(0,\sigma^{2}\mathbb{I})$ and $(-B,...,-B+1,...,-B)+\mathcal{N}(0,\sigma^{2}\mathbb{I})$ for $D$ and $D^{\prime}$, respectively. Since the inputs are shuffled, the “$-B+2$” and “$-B+1$” values can appear in any batch.
@@
Unlike DP-SGD, BGM is a hypothetical algorithm that is not already associated with common “data distributions” (e.g., training datasets). Therefore, when auditing BGM using the Distinguishability Game described earlier, we take the data distribution to be the “pathological” distribution, which always outputs ‘-1.’ Then, the adversary always chooses ‘+1’ as the target record $z_{T}$.
4.1 Single Epoch
@@
To audit a single epoch, an adversary has to distinguish between the outputs $(\tilde{g}_{1},...,\tilde{g}_{T})\sim\text{BGM}(D)$ and $(\tilde{g}^{\prime}_{1},...,\tilde{g}^{\prime}_{T})\sim\text{BGM}(D^{\prime})$. As mentioned, prior DP-SGD audits typically compute a natural “score” from each thresholded observation. However, in the context of shuffling, there is no known score to distinguish between the outputs.
@@
In this work, we draw from the Neyman-Pearson [46] lemma, which states that the optimal way to distinguish between two distributions is by thresholding the output of the likelihood ratio function. Specifically, the likelihood ratio function computes the ratio of probabilities that the outputs are from $\text{BGM}(D)$ or $\text{BGM}(D^{\prime})$. Here, the probability that $(\tilde{g}_{1},...,\tilde{g}_{T})$ is from $\text{BGM}(D)$ (or $\text{BGM}(D^{\prime})$) can be split into $T$ cases depending on which batch the target record $+1$ (or $\bot$) appears in. To that end, we calculate the likelihood ratio for the adjacent datasets $D=(+1,-1,...,-1)$ and $D^{\prime}=(\bot,-1,...,-1)$ as follows (we let $\tilde{\mathbf{g}}=(\tilde{g}_{1},...,\tilde{g}_{T})$):
$\small\begin{split}\Lambda(\tilde{\mathbf{g}})=&\frac{\Pr[\tilde{\mathbf{g}}|\text{BGM}(D)]}{\Pr[\tilde{\mathbf{g}}|\text{BGM}(D^{\prime})]}=\frac{\sum_{t}T^{-1}\Pr[\tilde{\mathbf{g}}|\text{BGM}(D)\land+1\text{ is in batch }t]}{\sum_{t}T^{-1}\Pr[\tilde{\mathbf{g}}|\text{BGM}(D^{\prime})\land\bot\text{ is in batch }t]}=\\ =&\frac{\sum_{t}\Pr[\tilde{g}_{t}|\mathcal{N}(-B+2,\sigma^{2})]\prod_{t^{\prime}\neq t}\Pr[\tilde{g}_{t^{\prime}}|\mathcal{N}(-B,\sigma^{2})]}{\sum_{t}\Pr[\tilde{g}_{t}|\mathcal{N}(-B+1,\sigma^{2})]\prod_{t^{\prime}\neq t}\Pr[\tilde{g}_{t^{\prime}}|\mathcal{N}(-B,\sigma^{2})]}\end{split}$
@@
4.2 Multiple Epochs
Although [16] only considers one set of batches (i.e., corresponding to a single “epoch” of training), we extend to multiple epochs as training private models typically involves training across multiple epochs. We rely on the fact that the batch sampling and aggregation are done independently across multiple epochs, thus, the likelihood across multiple epochs is the product of the likelihoods of each epoch, i.e., letting $\mathcal{G}=\begin{bmatrix}-\,\tilde{\mathbf{g}}^{1}-\\ \vdots\\ -\,\tilde{\mathbf{g}}^{E}-\end{bmatrix}$, we define the likelihood ratio as:
@@ 4.2 Multiple Epochs
Overall, our experimental evaluation (in Section 6.2) will show that auditing using likelihood ratios is very effective at estimating the privacy leakage from the Batched Gaussian Mechanism. In this idealized setting, we observe a substantial gap (up to 4$\times$) between the empirical privacy leakage observed and the theoretical guarantees produced by the Poisson sub-sampling analysis (recall that there are no known guarantees for shuffling). This prompts the need to investigate whether such a gap is also observable in more real-world settings using DP-SGD.
Our auditing techniques can be easily extended, as done above, by simply defining the appropriate likelihood ratio function. In the rest of the paper, we rely on this extensibility to audit the DP-SGD (Shuffle) implementation under varying threat models and audit implementations with variations in the shuffling procedure.

### v3

@@ 3.1 The Distinguishing Game
In Figure 2, we present the standard distinguishing game [42], involving an Adversary and a Challenger, adapted to audit the zero-out DP guarantees of a mechanism $\mathcal{M}$. The Challenger runs the mechanism on a randomly chosen dataset, and the Adversary aims to determine the dataset used by the Challenger based on the mechanism’s output.
In each game, the Adversary picks a single target record $z_{T}$ from the data domain and sends it to the Challenger. [^2] We are abusing notation by define $\mathcal{D}$ as both the data distribution and domain, although it is clear from the context. The latter constructs adjacent datasets $D_{0}$ and $D_{1}$ by appending the zero-out record $\bot$ and the target record $z_{T}$, respectively, to a randomly sampled dataset with n1n{-}1 records as done in prior work [42, 50, 2, 46]. Then, the Adversary is given access to the randomly sampled adjacent datasets $D_{0}$ and $D_{1}$. Next, the Challenger runs $\mathcal{M}$ on dataset $D_{b}$ for a random $b\in\{0,1\}$, and the Adversary wins if they correctly guess $\hat{b}=b$.
Note that $\mathcal{M}$ can be a complex algorithm with many implementation details that the theoretical privacy analysis does not rely on; thus, depending on the threat model considered and the adversarial capabilities, the adversary may be given access to specific internals of $\mathcal{M}$ as discussed in Section 5.1, along with $M$’s output. In theory, this can take any form, e.g., scalar $o\in\mathbb{R}$, vector $o\in\mathbb{R}^{d}$, or even have arbitrary domain $o\in\mathcal{Y}$. However, it can be difficult to consider and design decision functions around values in arbitrary domains. Hence, a common strategy [43, 31] is for the Adversary to define a distinguishing function $\mathcal{A}$ that assigns a scalar “score” to the output representing the Adversary’s confidence that the output is drawn from processing $D_{1}$. This score is then thresholded to produce a guess $\hat{b}=1$ if the Adversary determines that $o\sim\mathcal{M}(D_{1})$ and 0 otherwise.
@@ 3.2 Estimating $\varepsilon_{emp}$
We estimate the empirical privacy leakage $\varepsilon_{emp}$ by running the distinguishing game multiple times, where only a single record is inserted in each run. We compute the false positive rate (FPR) $\alpha$ and the false negative rate (FNR) $\beta$ across multiple games and derive (statistically valid) upper bounds $\overline{\alpha}$ and $\overline{\beta}$ using Clopper-Pearson confidence intervals (CIs) to quantify the confidence in our privacy loss estimation [42, 31]. We then calculate the empirical lower-bound privacy guarantee $\varepsilon_{emp}$ from $\overline{\alpha}$ and $\overline{\beta}$ using the ($\varepsilon,\delta$)-DP definition as follows.
While recent DP-SGD audits have presented “auditing in one run” methods [50, 41, 55], these do not yield sufficiently tight DP-SGD audits (even under powerful threat models), and thus we choose to audit using multiple runs. Moreover, prior work [42, 41, 55] has proposed alternative auditing techniques for DP-SGD (Poisson) using $f$-DP or PLD; however, since tighter $f$-DP guarantees for DP-SGD (Shuffle) are currently unknown [16], we stick to the ($\varepsilon,\delta$)-DP definition.
@@
$\begin{split}\mathcal{R}(\varepsilon,\delta)=\{(\alpha,\beta)|&\alpha+e^{\varepsilon}\beta\geq 1-\delta\land e^{\varepsilon}\alpha+\beta\geq 1-\delta~\land\\ &\alpha+e^{\varepsilon}\beta\leq e^{\varepsilon}+\delta\land e^{\varepsilon}\alpha+\beta\leq e^{\varepsilon}+\delta\}\end{split}$ (1)
@@
While Bayesian intervals [59] can improve the privacy estimation’s tightness, they are also more computationally expensive to derive and may not always be statistically sound [42]. Thus, we stick to Clopper-Pearson CIs. (In Section 6.2.1, we will also evaluate the impact of using CIs on the empirical privacy leakage estimation).
@@ 3.3 The Distinguishing Function
As mentioned, DP auditing involves an adversary distinguishing between observations from $\mathcal{M}(D)$ and $\mathcal{M}(D^{\prime})$. For a mechanism $\mathcal{M}:\mathcal{X}\rightarrow\mathcal{Y}$, this can be any function of the form $\mathcal{A}:\mathcal{Y}\rightarrow\{0,1\}$ where $\mathcal{A}(o)=1$ (or 0) represents that the adversary predicts that the observation $o$ is drawn from $\mathcal{M}(D)$ (or $\mathcal{M}(D^{\prime})$).
Scalar Score. In theory, $\mathcal{A}$ is equivalent to the rejection rule from the definition of $f$-DP, which represents the adversary performing a hypothesis test on whether to reject that $o\sim\mathcal{M}(D^{\prime})$. However, in previous work [32, 31, 43, 42], the adversary assigns a scalar “score” to each output, which is then thresholded to form the distinguishing function—e.g., Jagielski et al. [31] and Nasr et al. [42] use, respectively, the loss function and dot product. This is because the raw outputs of DP-SGD tend to be high-dimensional vectors where distances can be difficult to interpret, thus making distinguishing functions hard to design.
@@
Choosing the threshold. When auditing using scalar scores, choosing the threshold $\tau$ from an independent set of observations is the easiest method to ensure a technically valid lower bound for the empirical privacy leakage estimate $\varepsilon_{emp}$ derived this way. However, it is standard to report the maximum $\varepsilon_{emp}$ for the optimal threshold [3, 11, 31, 43, 42, 46, 59] and we do so too, making the process more efficient by first sorting the scalar scores. Here, we substantially improve the efficiency of this procedure by first sorting the scores in increasing order, as reported in Algorithm 1. This enables us to find the optimal threshold from, potentially, billions of observations without incurring a prohibitive computational cost. For simplicity, we abstract the details of choosing an appropriate threshold and designing a distinguishing function. We refer to the empirical estimation procedure at significance level $\alpha$ and privacy parameter $\delta$ from scores $\mathcal{S}$ and $\mathcal{S}^{\prime}$ derived from $\mathcal{M}(D)$ and $\mathcal{M}(D^{\prime})$, respectively as $\text{EstimateEps}(\mathcal{S},\mathcal{S}^{\prime},\alpha,\delta)$.
Algorithm 1 Estimating $\varepsilon_{emp}$ from scores
1:Scores from $\mathcal{M}(D)$, $\mathcal{S}$. Scores from $\mathcal{M}(D^{\prime})$, $\mathcal{S}^{\prime}$. Significance level, $\alpha$. Privacy parameter, $\delta$.
[…]
7:for $\tau\in\text{sort\_increasing}(\mathcal{S}\cup\mathcal{S}^{\prime})$ do
[…]
9:  if $\tau\in S$ then
10:    $\text{FNR}\leftarrow\text{FNR}+1$
11:  else
12:    $\text{FPR}\leftarrow\text{FPR}-1$
13:  $\triangleright$ Calculate confidence intervals for FPR and FNR.
14:  $\overline{\text{FPR}}\leftarrow\text{Clopper-Pearson}(FPR,R,\alpha)$
15:  $\overline{\text{FNR}}\leftarrow\text{Clopper-Pearson}(FNR,R,\alpha)$
16:  $\triangleright$ Estimate $\varepsilon_{emp}$ from Equation 2
17:  $\varepsilon_{emp}[\tau]=\max\left\{\ln\left(\frac{1-\overline{\text{FPR}}-\delta}{\overline{\text{FNR}}}\right),\ln\left(\frac{1-\overline{\text{FNR}}-\delta}{\overline{\text{FPR}}}\right),0\right\}$
18:return $\max_{\tau}\varepsilon_{emp}[\tau]$
4 Auditing the Batched Gaussian Mechanism
Before auditing DP-SGD (Shuffle), we first turn to the Batched Gaussian Mechanism (BGM) (see Algorithm 5 in Appendix A), a heavily simplified version of DP-SGD adapted from [16]. We do so to develop principled, (close to) tight auditing techniques for shuffling (under an idealized setting).
@@ 4 Auditing the Batched Gaussian Mechanism
Like DP-SGD, BGM also proceeds iteratively, sampling batches according to some batch sampler $\mathcal{B}$. However, instead of calculating gradients, the inputs are simply aggregated together with noise at each batch, and the noisy aggregated values are released for each batch across all epochs. Moreover, unlike DP-SGD, it is non-adaptive, i.e., the outputs of previous iterations do not affect the current one.
@@
While BGM can be instantiated with any batch sampler, we use the shuffle batch sampler, which is commonly used in ML training. We bound the inputs, i.e. $\forall i\;x_{i}\in[-1,+1]$, so that the mechanism satisfies $(\varepsilon,\delta)$-DP. When the number of epochs is 1, Chua et al. [16] use the adjacent datasets D=(+1,1,,1)D=(+1,-1,...,-1) and $D^{\prime}=(\bot,-1,...,-1)$, setting $\bot=0$ in their lower bound privacy analysis. Specifically, they conjecture, but do not prove, that these are the worst-case adjacent datasets for the shuffle batch sampler. Note that this is different from the proven worst-case datasets D=(+1,0,,0)D=(+1,0,...,0) and $D^{\prime}=(\bot,0,...,0)$ for the Poisson batch sampler [60]. Specifically, in the shuffle setting, the differing sample in the conjectured worst-case datasets is equal in magnitude but opposite in direction to the other samples, unlike in the Poisson setting. Finally, in a single epoch and for batch size $B$, the outputs of the mechanism on the adjacent datasets are $(-B,...,-B+2,...,-B)+\mathcal{N}(0,\sigma^{2}\mathbb{I})$ and $(-B,...,-B+1,...,-B)+\mathcal{N}(0,\sigma^{2}\mathbb{I})$ for $D$ and $D^{\prime}$, respectively. Since the inputs are shuffled, the “B+2-B+2” and “B+1-B+1” values can appear in any batch.
@@
We stress that BGM is a “hypothetical” algorithm, i.e., it is not used/associated with common data distributions or training datasets. Therefore, when auditing it using the Distinguishability Game, we take the data distribution to be the “pathological” distribution, which always outputs ‘-1.’ Then, the adversary always chooses ‘+1’ as the target record $z_{T}$.
4.1 From Single to Multiple Epochs
@@
Single Epoch. To audit a single epoch, an adversary has to distinguish between $(\tilde{g}_{1},...,\tilde{g}_{T})\sim\text{BGM}(D)$ and $(\tilde{g}^{\prime}_{1},...,\tilde{g}^{\prime}_{T})\sim\text{BGM}(D^{\prime})$. As mentioned, prior DP-SGD audits [32, 31, 43, 42] typically compute a natural “score” from each thresholded observation. However, in the context of shuffling, there is no known score to distinguish between the outputs.
@@
In this work, we draw from the Neyman-Pearson [44] lemma, which states that the optimal way to distinguish between two distributions is by thresholding the output of the likelihood ratio function. Specifically, the likelihood ratio function computes the ratio of probabilities that the outputs are from $\text{BGM}(D)$ or $\text{BGM}(D^{\prime})$. Here, the probability that $(\tilde{g}_{1},...,\tilde{g}_{T})$ is from $\text{BGM}(D)$ (or $\text{BGM}(D^{\prime})$) can be split into $T$ cases depending on which batch the target record +1+1 (or $\bot$) appears in. To that end, we calculate the likelihood ratio for the adjacent datasets D=(+1,1,,1)D=(+1,-1,...,-1) and $D^{\prime}=(\bot,-1,...,-1)$ as follows (we let $\tilde{\mathbf{g}}=(\tilde{g}_{1},...,\tilde{g}_{T})$):
[…]
@@
Multiple Epochs. Although [16] only considers one set of batches (i.e., corresponding to a single “epoch” of training), we extend to multiple epochs as training private models typically involves multiple epochs. We rely on the fact that the batch sampling and aggregation are done independently across multiple epochs; thus, the likelihood across multiple epochs is the product of the likelihoods of each epoch, i.e., letting $\mathcal{G}=\begin{bmatrix}-\,\tilde{\mathbf{g}}^{1}-\\ \vdots\\ -\,\tilde{\mathbf{g}}^{E}-\end{bmatrix}$, we define the likelihood ratio as:
@@
[…]
Figure 3: Auditing BGM for various number of steps $T$ and noise multipliers $\sigma$. Theoretical $\varepsilon$ (Poisson) represents the theoretical guarantees if Poisson subsampling had been used, while Theoretical $\varepsilon$ (Shuffle, lower bound) denotes the lower bounds from [16].
4.2 Experiments
4.2.1 Varying Batch Size and Noise Multiplier
We start by fixing both the number of epochs and the batch size ($B$) to 1. We vary the number of steps for a single epoch, $T$, varying the dataset size, and can then compute the theoretical Poisson analysis $\varepsilon$ using the privacy loss random variable (PRV) accountant for a given $T$ and noise multiplier $\sigma$. Furthermore, to assess computational complexity, we audit each setting over a varying number of game runs, each yielding a single observation. For context, we also consider the lower bound theoretical $\varepsilon$ previously calculated by Chua et al. [16].
Since the privacy loss distribution (PLD) used in [42] is not currently known for DP-SGD (Shuffle), recall that we are auditing using the $(\varepsilon,\delta)$-DP definition; as a result, we require a large number of observations for the $\varepsilon_{emp}$ estimates to converge, especially when $\sigma$ is small. For $\sigma\geq 1.0$, at $T=100$, $\varepsilon_{emp}$ converges after $10^{8}$ observations, but for $\sigma=0.5$, it requires more than $10^{9}$ observations (we do not explore beyond $10^{9}$ due to computational constraints).
As expected, given enough observations, our audits match but do not exceed the theoretical lower bounds calculated by Chua et al. [16]. This not only confirms that our auditing procedure is effective but also that it might be possible to convert the lower-bound privacy analysis from [16] into an upper-bound privacy guarantee. However, proving upper-bound privacy guarantees might not be trivial and is beyond our scope—thus, we leave it to future work.
4.2.2 Exploring parameter settings
Next, we audit using parameters from state-of-the-art differentially private models. Specifically, we audit BGM in the settings used in [19] and [39] to train private image classification models and large language models, respectively [^3] We audit the different parameter settings using our own shuffling implementation instead of auditing existing implementations both for computational efficiency and to freely vary the hyperparameters (e.g., batch size).. Note that De et al. [19] explicitly state they train their models with shuffling but report DP guarantees as though Poisson sampling was used. Whereas, Li et al. [39] do not mention using shuffling, but a review of their codebase shows they use the default Pytorch implementation of DataLoader, which shuffles the dataset.
[…]
Overall, these experiments indicate that, under the Batched Gaussian Mechanism’s ideal conditions, training with shuffling instead of Poisson subsampling incurs substantial privacy leakage, also in settings commonly used for training state-of-the-art models. This further motivates assessing the validity of the DP guarantees reported by state-of-the-art models.
Paper Dataset Size $\sigma$ $T$ $E$ $\varepsilon$ $\varepsilon_{emp}$ Gap
[…]
Table 1: Parameter settings used in prior work where the empirical privacy leakage ($\varepsilon_{emp}$) observed from shuffling is appreciably larger than the theoretical $\varepsilon$ from Poisson sampling.
4.3 Takeaways
As mentioned earlier, we experiment with Batched Gaussian Mechanism and study the impact of various parameter settings (noise scale, batch size, and number of epochs) on the empirical privacy leakage observed. Our experiments show that the empirical privacy leakage from shuffling can be substantially larger than the theoretical guarantees given by the Poisson subsampling analysis.
Crucially, this gap is prominent in the parameter settings used to train SOTA models in prior work [19, 39] as well. For instance, De et al. [19] reportedly use shuffling to fine-tune an NF-ResNet-50 model on the Places-365 dataset at a theoretical $\varepsilon=7.53$, but our audits show that the actual privacy leakage is almost double ($\varepsilon_{emp}=13.59$).